a semicircular driveway has a diameter of 180 ft. How long is the driveway?
Use 3.14 for Pie

Answer by jim_thompson5910(35256)

C = pi*d

C = 3.14*180

C = 565.2

So the circumference of the entire circle is 565.2 feet

Therefore, half of the circle is 565.2/2 = 282.6 feet (approximately)

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------